7. Find the value of the six trigonometric functions if the conditions provided hold. cos(20) = -and 180° SO S 270°
Answer #1

Solution : Cos (20) = - 1 2 cos (20) = 1-2 sin²o 1- 2 sin²o -1 j2 - 2 sim ²o - 1 - 1 2 = > 2 sino 1 1 V2 U 2 sin²o √2+1 V2 si=> costo 1 sin?o → + cos o sin²o 2 - =) Coso + 1 + 2 252 >> Cos o - + (1 + √2 252 =) cos o 1 + 2521-52 252 = cos o W VZ- 252- tano 11 1+52 22 12-1 252 - tan o 1+52 X √252 √252 vizi = tano [+ra √2-1 √2-1 (1+52) (12+1) (2-1) (2+1) (1+r2)2 (2)²-(1) ² 1Now, CSC (0) = 1 sin(0) 1 1+2 22 252 1+2 212 (1-12) (1+√2)(1-(2) 252 - 4 (1)? (2) ² 2((2-2) 1 - 2 2(12-2) -2(82-2) -√2(2-82)sec(o) 1 cos (6) 1 VE - 22 252 V - 2/2 (V2 + 1) V (12-))(2+1) 4 +22 2 - 1 2(2 +52) sec o √e (2+52) Now, 1 cot (0) = tan (0) 1cot(o) = 1 (3-252) (3 +252) ( 3 – 282) 3- 282 (3) -(252) 3 - 2 √2 9 - 8 3-22 - cot(o) 3- 252